Use x to represt my age in ears.
My brother is 5 years older than me.
The sum of our ages is 111 years.
Write an equation using the information given and work out my age.

Answer by ikleyn(52781)

.
Let x be your age.


Then your brothers age is (x+5) years, according to the condition.


The sum of your ages is 111 years


    x + (x+5) = 111.


It is the equation the problem asks to write.


Solve it


    2x + 5 = 111

    2x     = 111 - 5

    2x     = 106

     x     = 106/2 = 53.


Your age is 53 years.                        ANSWER


Your brother's age is  53 + 5 = 58 years.    ANSWER


CHECK.  53 + 58 = 111  years.    ! Correct !

Solved.